# Partial Functional Dynamic Backdoor Diffusion-based Causal Model

### v1

@@ 4 Bounding Counterfactual Error
In this section, we establish the theoretical guarantees for the counterfactual estimation accuracy of the PFD-BDCM framework. The primary contribution is the derivation of an error bound that formally links the reconstruction fidelity of the encoder-decoder architecture to the precision of its counterfactual predictions. Notably, these theoretical guarantees accommodate higher-dimensional settings through additional structural assumptions. Formal proofs are presented in Appendix B.
@@
$g$ $:\mathcal{X}\times\mathcal{X}_{\mathcal{B}}\to\mathcal{Z}\quad\text{(encoding function)},$
$h$ $:\mathcal{Z}\times\mathcal{X}_{\mathcal{B}}\to\mathcal{X}\quad\text{(decoding function)},$
where $\mathcal{Z}$ denotes the latent space. Within PFD-BDCM, $g$ and $h$ correspond to the $\operatorname{Enc}$ and $\operatorname{Dec}$ operators, respectively.
@@
These assumptions, while formal, are well-motivated in the context of causal inference and deep generative models. Assumption 1 ensures that the encoder learns a representation of the exogenous noise that is not contaminated by information from the backdoor variables. This is naturally satisfied in settings like additive noise models with $f_{ij}(X_{\mathcal{B}},U)=f_{ij}^{\prime}(X_{\mathcal{B}})+U$ where $X_{\mathcal{B}}$ and $U$ is independent. If the fitted model $\hat{f}_{ij}\equiv f_{ij}^{\prime}$, then $g(X,X_{\mathcal{B}})=U$. Assumption 2 is satisfied by major identifiable model classes-including additive noise, post-nonlinear, and heteroscedastic formulations (Strobl and Lasko,, 2023) while concurrently resolving symmetric noise ambiguities characteristic of observational data. This assumption further aligns with contemporary identifiability frameworks (Nasr-Esfahany and Kiciman,, 2023) and intrinsically precludes non-identifiable structural equations. The encoder invertibility condition (Assumption 3) is intrinsically satisfied by the bijective properties of deterministic diffusion architectures (Song et al.,, 2021), guaranteeing uniqueness in latent representations while preserving compatibility with standard implementations.
@@
This theorem provides the foundation for assessing the accuracy of counterfactuals given by the PFD-BDCM. It implies that the abduction step, $\operatorname{Enc}(X,X_{\mathcal{B}})$, correctly captures the essence of the unobserved confounder $u$ that generated the factual observation. We now explore the direct consequences of this result.
Perfect Estimation.
@@
This corollary formally establishes that minimizing the reconstruction loss during training directly optimizes the model for better counterfactual prediction.
The framework extends to the more general multivariate setting where $X\in\mathbb{R}^{d}$. This requires a stronger assumption on the encoder’s Jacobian to ensure that information is not lost in the higher-dimensional space.
Theorem 2.
Consider $X\in\mathcal{X}\subset\mathbb{R}^{d}$ and $U\in\mathbb{R}^{m}$ with $m\geq d$. Assume the following conditions:
  1. 1.
    The structural equation $X=f_{ij}(X_{\mathcal{B}},U)$ is bijective in $U$ for any fixed $X_{\mathcal{B}}$.
  2. 2.
    The encoder’s Jacobian, $\partial g/\partial X$, has full rank $d$ almost everywhere.
  3. 3.
    The decoder $h(Z,X_{\mathcal{B}})$ is $L_{h}$-Lipschitz with respect to its first argument $Z$.
  4. 4.
    The reconstruction error is bounded such that $\|h(g(X,X_{\mathcal{B}}),X_{\mathcal{B}})-X\|\leq\tau$.
Then for a counterfactual intervention $\mathrm{do}(X_{\mathcal{B}}:=\gamma)$, the estimation error is bounded by:
$\|\hat{x}^{\mathrm{CF}}-x^{\mathrm{CF}}\|\leq L_{h}\cdot\kappa_{g}\cdot\tau,$
where $\kappa_{g}=\sup\|(\partial g/\partial X)^{-1}\|_{2}$ is the bound for the inverse of the encoder’s Jacobian.
@@ 5 Experimental Evaluation
We empirically evaluated the efficacy of PFD-BDCM in addressing causal queries across synthetic and real-world datasets. All experiments were implemented in $\mathrm{Python}$ (version 3.9.0) on a Windows server equipped with an Intel i9-13900K processor, NVIDIA RTX 4090 24G GPU, and 128 GB RAM. To demonstrate that PFD-BDCM faithfully samples from the target interventional distribution, we designed scenarios where causal sufficiency was deliberately violated within partially functional structural models.
@@ 5.1 Simulation Study
Figure 2 depicted in Appendix C.1 presented the instantiated PFST-DSCM, where causal sufficiency was compromised. Consider $\{X_{28},X_{29}\}$ as unobserved explanatory variables and $X_{30}$ as an unobserved explained variable, we assumed exhibit pronounced spatial heterogeneity coupled with temporal dependence between $\{X_{28},X_{29}\}$ and $X_{30}$. Let $X_{1},X_{2},X_{3}$ represent endogenous cause variables; $X_{31},X_{32},X_{33}$ denote outcome variables; and $\boldsymbol{X}_{\mathcal{B}}=\{X_{22},\cdots,X_{27}\}$ constitute backdoor adjustment sets. For visual clarity, exogenous noise terms $\boldsymbol{U}$ were omitted.
The partial functional dynamic structural equations were defined as: $X_{ijk}=f_{ij}(X_{\mathcal{B}k},U_{k})$ (Eq. (23) in Appendix C.1). The structural equations governing the Partially Functional Dynamic Diffusion-based Causal Model(PFD-DCM) and PFD-BDCM were instantiated with additive noise models (ANM) (Peters et al.,, 2013) furnishing elementary baselines.
Our objective was to accurately sample from the post-interventional distribution $q(X_{k}|\operatorname{do}(X_{l}=\gamma_{l}))$, where $k\in\{31,32,33\}$ indexes outcomes and $l\in\{1,2,3\}$ indexes causes. During intervention, $X_{l}$ is fixed to $\gamma_{l}$, while root variables $X_{h}$ ($h=1,\cdots,21$) were sampled from their empirical marginals $E_{h}$. For outcome $X_{k}$, PFD-DCM employed $\operatorname{Dec}_{k}(Z_{k},\hat{X}_{h})$, whereas PFD-BDCM utilized $\operatorname{Dec}_{k}(Z_{k},\hat{X}_{\mathcal{B}_{k}},\hat{X}_{h})$, thereby leveraging backdoor adjustments. Comprehensive simulation setting was detailed in the Appendix C.1.
Table 1 summarizes aggregated performance metrics-observational (Obs.), interventional (Int.), and counterfactual (CF.)-averaged over nine independent random initializations. Comprehensive diagnostics (boxplots and kernel density estimates) were provided in Appendix C.1. PFD-DCM and PFD-BDCM achieved compelling statistical fidelity across all query types, evidenced by MMD and CRPS for observational/interventional query and low MSE for counterfactual query. Crucially, PFD-BDCM consistently outperformed PFD-DCM, attributable to its principled utilization of backdoor adjustments that mitigated information loss from critical latent nodes.
$J=6$ PFD-DCM PFD-BDCM
$n=30$ $3.925\pm 3.792$ $3.532\pm 3.920$
@@
(Obs.) $n=200$ $15.309$ $15.671$
$n=500$ $44.661$ $44.918$
$n=30$ $4.021\pm 3.749$ $4.298\pm 3.946$
@@
(CF.) $n=200$ $0.074\pm 0.020$ $0.598\pm 0.095$
$n=500$ $0.024\pm 0.006$ $0.579\pm 0.056$
$n=30$ $1.537$ $1.535$
@@
(CF.) $n=200$ $8.621$ $8.616$
$n=500$ $22.163$ $22.167$
[…]
5.2 Empirical Application
This investigation employed the PFD-BDCM framework to examine spatio-temporal dynamic structural causal relationships among air pollutant indicators and their determinants. Our analysis encompassed 30 provincial-level administrative divisions across Chinese mainland during the period January 2015 to December 2020. The study integrates China’s provincial CO2 emission inventories from the China Emission Accounts and Datasets (CEADs) (Shan et al.,, 2018, 2020; Guan et al.,, 2021; Xu et al.,, 2024) and emissions data for nine atmospheric pollutants from the Multi-scale Emission Inventory of China (MEIC) (Li et al.,, 2018; Zheng et al.,, 2018; Geng et al.,, 2024; Li et al.,, 2019) as response variables for air pollutant emissions.
Building upon prior research (Ghosh,, 2010; Jayanthakumaran et al.,, 2012; Ozcan,, 2013; Zhu et al.,, 2021) and incorporating domain-specific characteristics of regional emissions, we systematically collected foundational determinants across ten conceptual dimensions. The comprehensive dataset comprised 118 indicator variables, through collinearity diagnostics and random forest-based feature selection, we retained 49 statistically robust indicators for subsequent modeling (detailed indicators show in Appendix C.2 Table 6).
Complete experimental specifications and supplementary materials were documented in Appendix C.2, with observational query results presented in Table 2 .
Variable PFD-DCM PFD-BDCM
SO2 $0.495\pm 0.419$ $0.512\pm 0.476$
NOx $0.511\pm 0.449$ $0.473\pm 0.424$
CO $0.564\pm 0.531$ $0.524\pm 0.450$
VOC $0.489\pm 0.464$ $0.480\pm 0.403$
NH3 $0.519\pm 0.538$ $0.494\pm 0.446$
PM10 $0.487\pm 0.426$ $0.485\pm 0.457$
PM2.5 $0.400\pm 0.356$ $0.435\pm 0.400$
BC $0.417\pm 0.406$ $0.350\pm 0.339$
OC $0.487\pm 0.462$ $0.398\pm 0.413$
CO2 $0.336\pm 0.257$ $0.357\pm 0.304$
[…]
@@ 6 Concluding Remarks
This research introduces the Partial Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), a novel methodological framework crafted for robust causal inference amidst spatial heterogeneity, temporal dependencies, and unmeasured confounding. Our principal contributions are threefold.
Model Innovation: PFD-BDCM synergistically integrates functional basis expansions with diffusion-based causal modeling, facilitating simultaneous resolution of: i) Multi-resolution variables through partial functional representations; ii) Spatio-temporal dynamics via regionally parameterized structural equations; iii) Unmeasured confounder bias utilizing backdoor adjustment sets.
@@
While PFD-BDCM advances causal inference in complex settings, future work should address: i) Scalability enhancements for ultra-high-dimensional functional data via tensor decompositionsp; ii)Automated backdoor set identification through causal discovery algorithms; iii)Temporal graph neural network integration for non-stationary processes; iv) Real-time deployment in environmental policy decision support systems.
The proposed framework opens new avenues for causal inference in environmental science, epidemiology, and econometrics where functional data and unmeasured confounders are prevalent.

### v2

@@ 3 Counterfactual Error Bounds
In this section, we establish the theoretical guarantees for the counterfactual estimation accuracy of the PFD-BDCM framework. The primary contribution is the derivation of an error bound that formally links the reconstruction fidelity of the encoder-decoder architecture to the precision of its counterfactual predictions. Notably, these theoretical guarantees accommodate higher-dimensional settings through additional structural assumptions. Formal proofs are presented in Appendix C.
@@
$g:\mathcal{X}\times\mathcal{X}_{\mathcal{B}}\to\mathcal{Z}\quad\text{(encoding function)};\quad h:\mathcal{Z}\times\mathcal{X}_{\mathcal{B}}\to\mathcal{X}\quad\text{(decoding function)},$
where $\mathcal{Z}$ denotes the latent space. Within PFD-BDCM, $g$ and $h$ correspond to the $\mathrm{Enc}$ and $\mathrm{Dec}$ operators, respectively.
@@
These assumptions, while formal, are well-motivated in the context of causal inference and deep generative models. Assumption 1 ensures that the encoder learns a representation of the exogenous noise that is not contaminated by information from the backdoor variables. This is naturally satisfied in settings like additive noise models with $f_{ij}(X_{\mathcal{B}},U)=f_{ij}^{\prime}(X_{\mathcal{B}})+U$ where $X_{\mathcal{B}}$ and $U$ is independent. If the fitted model $\hat{f}_{ij}\equiv f_{ij}^{\prime}$, then $g(X,X_{\mathcal{B}})=U$. Assumption 2 is satisfied by major identifiable model classes, including additive noise, post-nonlinear, and heteroscedastic formulations (Strobl and Lasko,, 2023) while concurrently resolving symmetric noise ambiguities characteristic of observational data. This assumption further aligns with contemporary identifiability frameworks (Nasr-Esfahany and Kiciman,, 2023) and intrinsically precludes non-identifiable structural equations. The encoder invertibility condition (Assumption 3) is intrinsically satisfied by the bijective properties of deterministic diffusion architectures (Song et al.,, 2021), guaranteeing uniqueness in latent representations while preserving compatibility with standard implementations.
@@
Theorem 1 provides the foundation for assessing the accuracy of counterfactuals given by the PFD-BDCM. It implies that the abduction step, $\mathrm{Enc}(X,X_{\mathcal{B}})$, correctly captures the essence of the unobserved confounder $u$ that generated the factual observation. We now explore the direct consequences of this result.
@@
This corollary formally establishes that minimizing the reconstruction loss during training directly optimizes the model for better counterfactual prediction. The framework extends to the more general multivariate setting where $X\in\mathbb{R}^{d}$ (Theorem 2 ). This requires a stronger assumption on the encoder’s Jacobian to ensure that information is not lost in the higher-dimensional space. We present it in Appendix C.
@@ 4 Experimental Evaluation
We empirically evaluated the efficacy of PFD-BDCM in addressing causal queries across synthetic and real-world datasets. To demonstrate that PFD-BDCM faithfully samples from the target interventional distribution, we designed scenarios where causal sufficiency was deliberately violated within Partially functional structural models.
@@ 4.1 Simulation Study
Figure 2: PFST-DSCM with 33 exogenous and endogenous nodes (where nodes $X_{28}$, $X_{29}$ and $X_{30}$ are unmeasured confounders with spatial heterogeneity and temporal dependencies, $Y_{1}(t),Y_{2}(t),Y_{3}(t)$ are functional nodes, and $X_{4},\cdots,X_{21}$ are the corresponding base expansion nodes)
Figure 2 presented the instantiated PFST-DSCM, where causal sufficiency was compromised. Consider $\{X_{28},X_{29}\}$ as unobserved explanatory variables and $X_{30}$ as an unobserved explained variable, we assumed exhibit pronounced spatial heterogeneity coupled with temporal dependence between $\{X_{28},X_{29}\}$ and $X_{30}$. Let $X_{1},X_{2},X_{3}$ represent endogenous cause variables; $X_{31},X_{32},X_{33}$ denote outcome variables; and $\boldsymbol{X}_{\mathcal{B}}=\{X_{22},\cdots,X_{27}\}$ constitute backdoor adjustment sets. For visual clarity, exogenous noise terms $\boldsymbol{U}$ were omitted.
The partially functional dynamic structural equations were defined as: $X_{ijk}=f_{ij}(X_{\mathcal{B}k},U_{k})$ (Eq. (37) in Appendix D ). The structural equations governing the Partially Functional Dynamic Diffusion-based Causal Model (PFD-DCM) and PFD-BDCM were instantiated with additive noise models (ANM) (Peters et al.,, 2013) furnishing elementary baselines.
Our objective was to accurately sample from the post-interventional distribution $q(X_{k}|\operatorname{do}(X_{l}=\gamma_{l}))$, where $k\in\{31,32,33\}$ indexes outcomes and $l\in\{1,2,3\}$ indexes causes. During intervention, $X_{l}$ is fixed to $\gamma_{l}$, while root variables $X_{h}$ ($h=1,\cdots,21$) were sampled from their empirical marginals $E_{h}$. For outcome $X_{k}$, PFD-DCM (DCM) employed $\mathrm{Dec}_{k}(Z_{k},\hat{X}_{h})$, whereas PFD-BDCM (BDCM) utilized $\mathrm{Dec}_{k}(Z_{k},\hat{X}_{\mathcal{B}_{k}},\hat{X}_{h})$, thereby leveraging backdoor adjustments. Comprehensive simulation setting was detailed in the Appendix D.
@@
$n=200$ $15.671$ $15.309$ $15.665$ $\mathbf{15.306}$
$n=30$ $\mathbf{3.580\pm 3.013}$ $4.282\pm 3.952$ $3.990\pm 3.962$ $3.922\pm 3.247$
$\downarrow$ MMD (Int.) $n=80$ $\mathbf{1.408\pm 1.204}$ $1.511\pm 1.268$ $2.009\pm 2.052$ $2.079\pm 2.003$
$n=200$ $0.595\pm 0.504$ $\mathbf{0.592\pm 0.582}$ $2.803\pm 3.620$ $3.187\pm 3.163$
$n=30$ $2.291$ $2.171$ $2.286$ $\mathbf{2.165}$
@@
$n=200$ $8.616$ $8.621$ $8.173$ $\mathbf{8.023}$
Table 1 summarizes aggregated performance metrics-observational (Obs.), interventional (Int.), and counterfactual (CF.)-averaged over nine independent random initializations. Comprehensive diagnostics (boxplots (Fig. 4 and 5 ) and kernel density estimates(Fig. 3 )) were provided in Appendix D . PFD-DCM and PFD-BDCM achieved compelling statistical fidelity across all query types, evidenced by MMD for observational/interventional query and MSE for counterfactual query. PFD-BDCM consistently outperforms baselines, demonstrating superior MMD metrics in observational queries and enhanced stability in counterfactual queries under varying data scales. This advantage stems from its principled integration of backdoor adjustment with spatiotemporal modeling of unobserved confounders, which mitigates information loss in latent nodes and corrects confounding-induced biases, significantly improving causal query performance.
4.2 Empirical Application
This investigation employed the PFD-BDCM framework to examine spatio-temporal dynamic structural causal relationships among air pollutant indicators and their determinants. Our analysis encompassed 30 provincial-level administrative divisions across Chinese mainland during the period January 2015 to December 2020. The study integrates China’s provincial CO2 emission inventories from the China Emission Accounts and Datasets (CEADs) (Guan et al.,, 2021; Xu et al.,, 2024) and emissions data for nine atmospheric pollutants from the Multi-scale Emission Inventory of China (MEIC) (Li et al.,, 2019; Geng et al.,, 2024) as response variables for air pollutant emissions.
Building upon prior research (Ozcan,, 2013; Zhu et al.,, 2021) and incorporating domain-specific characteristics of regional emissions, we systematically collected foundational determinants across ten conceptual dimensions. The comprehensive dataset comprised 118 indicator variables, through collinearity diagnostics and random forest-based feature selection, we retained 49 statistically robust indicators for subsequent modeling (detailed indicators shown in Appendix D.2 Table 8.Complete experimental specifications and supplementary materials were documented in Appendix D.2 , with observational query results presented in Table 2.
@@ 4.2 Empirical Application
Variable PFD-DCM PFD-BDCM Variable PFD-DCM PFD-BDCM
SO2 $\mathbf{0.495\pm 0.419}$ $0.512\pm 0.476$ PM10 $0.487\pm 0.426$ $\mathbf{0.485\pm 0.457}$
NOx $0.511\pm 0.449$ $\mathbf{0.473\pm 0.424}$ PM2.5 $\mathbf{0.400\pm 0.356}$ $0.435\pm 0.400$
CO $0.564\pm 0.531$ $\mathbf{0.524\pm 0.450}$ BC $0.417\pm 0.406$ $\mathbf{0.350\pm 0.339}$
VOC $0.489\pm 0.464$ $\mathbf{0.480\pm 0.403}$ OC $0.487\pm 0.462$ $\mathbf{0.398\pm 0.413}$
NH3 $0.519\pm 0.538$ $\mathbf{0.494\pm 0.446}$ CO2 $\mathbf{0.336\pm 0.257}$ $0.357\pm 0.304$
@@ 5 Concluding Remarks
We propose the Partially Functional Dynamic Backdoor Diffusion-based Causal Model (PFD-BDCM), a methodological framework crafted for robust causal inference amidst spatial heterogeneity, temporal dependencies, and unmeasured confounding. Our contributions are threefold.
Model Innovation: PFD-BDCM synergistically integrates functional basis expansions with diffusion-based causal modeling, facilitating simultaneous resolution of: i) Multi-resolution variables through partially functional representations; ii) Spatio-temporal dynamics via regionally parameterized structural equations; iii) Unmeasured confounder bias utilizing backdoor adjustment sets.
@@
While PFD-BDCM advances causal inference in complex settings, future work should address: i) Scalability enhancements for ultra-high-dimensional functional data via tensor decomposition; ii) Automated backdoor set identification through causal discovery algorithms; iii) Temporal graph neural network integration for non-stationary processes; iv) Real-time deployment in environmental policy decision support systems.
The proposed framework opens new avenues for causal inference in environmental science, epidemiology, and econometrics, where functional data and unmeasured confounders are prevalent.